OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-4,2) and y-intercept (0,-30). Write your answer in this form: (x1, y1),(x2, y2). If necessary, round to the nearest hundredth

OpenStudy (zepp):

First find the equation of this parabola, you have the vertex and the y-intercept: \(y=a(x-h)^2+k\\ -30=a(0-(-4))^2+2\\ -30=a(4)^2+2\\ 16a+2=-30\\ -30-2=16a\\ -32=16a\\ a=-32/16\\ a=-2\) So the equation of this parabola would be \(y=-2(x+4)^2+2\) Now to find the x-intercepts, replace y by 0 and solve for x.

OpenStudy (anonymous):

I got x=2...

OpenStudy (zepp):

There's two of them. Remember, a square root always gives 2 answers, e.g.: \(\sqrt{25}=-5 ~and~ 5\)

OpenStudy (anonymous):

I put it in as (-5,0), (5,0) It was wrong..

OpenStudy (zepp):

It's an example D:

OpenStudy (anonymous):

oh... Well the one I'm on now has a vertex of (-7,45) nd y-intercept of (0,-200)

OpenStudy (zepp):

\( 0=−2(x+4)^2+2\\ -2=-2(x+4)^2\\ 1=(x+4)^2\\ \pm \sqrt{1}=x+4\) So we have -1 = x+4; x=-1-4=-5 and 1=x+4; x=1-4=-3 Roots at (-5,0) and (-3,0)

OpenStudy (anonymous):

That was wrong too.. I think I might be typing it wrong...

OpenStudy (anonymous):

Vertex of (2, 13) and y-intercept of (0, 5)

OpenStudy (zepp):

That one, is the previous question...

OpenStudy (anonymous):

...I might throw my computer.. D: I'm dumb lgogbe

OpenStudy (anonymous):

I got it now, thanks!

OpenStudy (zepp):

You are welcome :)

OpenStudy (anonymous):

Can you help me with another one? It's my last one an I need to get it right. Find the x-intercepts of the parabola with vertex (4,-1) and y-intercept (0,15). Answer in this form: (x1,y1),(x2,y2)